### v1

# Enhancing Reasoning Skills in Small Persian Medical Language Models Can Outperform Large-Scale Data Training

Abstract
Enhancing reasoning capabilities in small language models is critical for specialized applications such as medical question answering, particularly in underrepresented languages like Persian. In this study, we employ Reinforcement Learning with AI Feedback (RLAIF) and Direct preference optimization (DPO) to improve the reasoning skills of a general-purpose Persian language model. To achieve this, we translated a multiple-choice medical question-answering dataset into Persian and used RLAIF to generate rejected-preferred answer pairs, which are essential for DPO training. By prompting both teacher and student models to produce Chain-of-Thought (CoT) reasoning responses, we compiled a dataset containing correct and incorrect reasoning trajectories. This dataset, comprising 2 million tokens in preferred answers and 2.5 million tokens in rejected ones, was used to train a baseline model, significantly enhancing its medical reasoning capabilities in Persian. Remarkably, the resulting model outperformed its predecessor, gaokerena-V, which was trained on approximately 57 million tokens, despite leveraging a much smaller dataset. These results highlight the efficiency and effectiveness of reasoning-focused training approaches in developing domain-specific language models with limited data availability.
I Introduction
In 2019, at the NeurIPS conference, Yoshua Bengio highlighted a critical limitation of current deep learning systems: their inability to perform tasks requiring robust reasoning skills [1]. While these systems excel at intuitive, perception-driven tasks, they struggle with reasoning-based challenges that demand deeper cognitive processing. This observation echoes Daniel Kahneman’s distinction between “fast” and “slow” thinking in cognitive science [2]. Fast, intuitive thinking aligns with the strengths of current AI systems, whereas slow, deliberate reasoning remains a significant weakness.
One proposed solution to this deficiency is the revival of symbolic AI methods through neurosymbolic approaches. However, Bengio cautioned against such methods due to scalability issues. Instead, he urged the development of new architectures capable of out-of-distribution generalization—a core requirement for genuine reasoning.
The Transformer architecture [3], which underpins most modern language models, suffers from the same deficiency. Transformers often make “stupid mistakes” when faced with reasoning-intensive problems, prompting researchers to enlarge models and datasets in an attempt to achieve better results across tasks. This approach arguably simulates System 2 reasoning within a fundamentally System 1 framework.
This limitation becomes even more pronounced in smaller language models, where restricted capacity and limited data exacerbate reasoning challenges. In domains such as medicine—where reasoning, interpretation, and decision-making are essential—these weaknesses become especially consequential.
Recent efforts (including ours) to enhance reasoning have largely focused on prompting or fine-tuning methods that enable models to imitate reasoning behavior rather than truly engage in it—a phenomenon we refer to as the “illusion of thinking” [4]. This means that such approaches provide only an approximation of genuine cognitive reasoning. Nevertheless, in the absence of architectures with intrinsic reasoning capabilities, these efforts represent necessary steps toward developing AI systems with more robust reasoning abilities.
The present work aims to address this challenge by enhancing the reasoning abilities of a small Persian language model, aya-expanse-8b [5]. We first prompt the model to generate Chain-of-Thought (CoT) trajectories for multiple-choice medical questions. These responses are then reviewed and corrected using DeepSeek-R, a model adept at reasoning, to rectify errors. This process constitutes Reinforcement Learning with AI Feedback (RLAIF) [6]. By pairing incorrect and corrected CoT trajectories, we subsequently apply Direct Preference Optimization (DPO) [7] to enhance the reasoning capabilities of the baseline model.
We call this improved model gaokerena-R. While its predecessor, gaokerena-V [8], demonstrated superior medical knowledge in straightforward prompting tasks, gaokerena-R outperformed it when evaluated using Chain-of-Thought (CoT) prompting [9]. Notably, gaokerena-R was trained on a much smaller dataset, focusing on reasoning enhancement rather than data scale, whereas gaokerena-V relied on extensive training on a large medical corpus and the MF3QA dataset [8].
Our findings highlight an important insight: in low-resource domains such as Persian medical AI, strengthening reasoning capabilities can be more effective than training on vast datasets. This underscores the importance of developing reasoning-oriented approaches, particularly in languages and fields with limited data availability. By improving reasoning abilities in smaller models, we can make meaningful progress toward practical, resource-efficient AI systems.
II Related Work
To the best of our knowledge, no prior work has focused on developing Persian medical reasoning language models. Existing Persian medical models, including our previous gaokerena-V, mainly address knowledge representation and general language understanding with limited attention to reasoning.
L. Pan et al. reviewed several valuable studies on enhancing reasoning capabilities in language models. [10] Accordingly, we focus here on English-language works that offer methodological insights into generating and integrating reasoning data for improving medical language models.
II-A Related Work In Medical Domain
One of the notable efforts in developing small-scale medical reasoning language models is MedSSS [11]. This framework focuses on fine-graining the intermediate reasoning steps to enhance the reasoning ability of medical models. To accomplish this, the authors applied the Monte Carlo Tree Search (MCTS) [12] algorithm on medical multiple-choice question-answering datasets to synthesize structured reasoning trajectories using a policy model. Based on this approach, they constructed three datasets supporting different stages of training: a dataset for Supervised Fine-Tuning (SFT) of the policy model, a rejected–preferred answer dataset for Direct Preference Optimization (DPO) [13] ,and a soft dual-side label dataset for fine-tuning the Process Reward Model (PRM).Finally, the trained Policy Model was used as the primary reasoning policy, while the trained Process Reward Model acted as a decoding guide to refine and evaluate the reasoning process during generation.
Another significant contribution in this field is MedReason [14]. In this work, the authors leveraged a structured medical knowledge graph to transform conventional question–answer (Q&A) pairs into detailed medical reasoning trajectories. Each trajectory represented the step-by-step logical pathway connecting a clinical question to its correct answer, grounded in medical knowledge relationships such as symptoms, diagnoses, treatments, and physiological mechanisms. By constructing a reasoning-enriched dataset in this manner, the authors were able to fine-tune a baseline language model on these structured reasoning examples. This approach led to a marked improvement in the model’s ability to perform reasoning-intensive tasks within the medical domain, demonstrating the effectiveness of incorporating knowledge graph–based reasoning supervision into the training process. The findings from MedReason further emphasize the importance of structured reasoning representations as a means of improving the interpretability and analytical depth of medical language models.
Another important advancement in the development of medical reasoning models is HuatuoGPT-o1 [15]. In this work, the authors introduced a verification-guided reasoning framework designed to improve the logical consistency and accuracy of generated reasoning trajectories. Specifically, they employed a verifier model to assess and guide the policy model during reasoning generation, ensuring that each reasoning path adhered closely to factual correctness and medical plausibility. By filtering and refining reasoning trajectories through this verification process, they created a high-quality dataset composed of verified reasoning sequences. The authors then leveraged both supervised fine-tuning and reinforcement learning techniques to train their baseline language model using these verified data. This dual training approach allowed the model to not only imitate correct reasoning behaviors but also internalize reasoning principles through reward-driven optimization. As a result, HuatuoGPT-o1 demonstrated significant improvements in reasoning accuracy and reliability across a variety of medical question-answering and diagnostic tasks, highlighting the potential of verifier-guided learning frameworks in advancing medical reasoning language models.
II-B Related Work In Other Domain
Perhaps the most influential recent work in the broader field of AI reasoning is DeepSeek-R [16]. Building upon the DeepSeek-V3-Base model, the authors introduced a reinforcement learning framework based on Gradient Regularized Policy Optimization (GRPO) [17] to explicitly enhance the model’s reasoning capabilities. In this setup, the reinforcement learning process was guided by a reward function specifically designed to evaluate and maximize reasoning performance. Through this training paradigm, DeepSeek-R demonstrated remarkable improvements in logical reasoning and problem-solving accuracy across various benchmarks. However, the application of reinforcement learning also introduced several side effects. While reasoning performance improved substantially, the model’s readability and linguistic coherence deteriorated, and instances of language mixing became more frequent. To address these issues, the authors incorporated a small amount of cold-start supervised data and adopted a multi-stage training pipeline. This additional fine-tuning phase helped restore natural language fluency and readability while retaining the strong reasoning skills acquired through reinforcement learning. The resulting model, DeepSeek-R, thus represents a critical step forward in reasoning-oriented AI, demonstrating that reinforcement learning can significantly enhance reasoning ability—provided it is balanced with targeted fine-tuning to maintain linguistic quality.
Another notable contribution in the area of reasoning enhancement is Thought Preference Optimization (TPO) [18]. In this work, the authors proposed a preference-based framework for improving reasoning quality in language models. Given a question, the model first generates multiple candidate reasoning trajectories. These responses are then evaluated by a judge model, which identifies the best and worst samples based on reasoning correctness and coherence. The collected best–worst pairs are subsequently used to train the baseline model through Direct Preference Optimization (DPO), encouraging it to prefer higher-quality reasoning paths. This approach effectively aligns the model’s reasoning process with human-like evaluative feedback, demonstrating that reasoning quality can be substantially improved through structured preference optimization rather than relying solely on scale or supervised data.
Another relevant study was conducted by N. Ho et al. [19], who explored the transfer of reasoning capabilities from large language models to smaller ones. In their approach, a smaller model was fine-tuned using data generated by a larger model that exhibited stronger reasoning performance. The larger model produced reasoning trajectories and question–answer pairs that served as high-quality supervision signals for the smaller model. Through this distillation process, the smaller model effectively learned reasoning strategies and problem-solving patterns from its larger counterpart, achieving competitive reasoning performance with significantly reduced computational cost. This work demonstrates that reasoning ability can be efficiently transferred across models of different scales through targeted fine-tuning on reasoning-oriented synthetic data.
III Data
We required a Persian medical multiple-choice question answering dataset to evaluate our proposed methods. In the absence of such a dataset, we used DeepSeek-V3 [20] , a cost-effective large language model, to translate a small portion of the MedMCQA dataset from English to Persian. We randomly selected these questions from MedMCQA [21] to maintain topic diversity in the resulting dataset. To verify the quality of the translations, we prompted  [^1] Prompts are available at github.com/Mehrdadghassabi/Gaokerena-R/blob/main/dataset/judgement.ipynb two referees (grok-3-mini [22] and gpt-4.1-mini [23] ) to evaluate each translation. A translation was considered verified only if the both referees assigned it a score of 5 out of 5. This process resulted in approximately 11,000 Persian medical multiple-choice questions.
III-A PersianMedQA
Shortly after completing our translation process, a new Persian medical multiple-choice question answering dataset called PersianMedQA was introduced [24] . We randomly selected a small portion of its training data and appended it to our dataset to include additional data from a different source.
IV Proposed Methods
We proposed two methods. The first framework, depicted in 1, utilizes a teacher model ( we used DeepSeek-R), which is itself a reasoning language model, to correct the student model’s medical reasoning mistakes in its response to the given multiple choice question. (if the student model chose the correct option we move to the next question in our dataset) When provided with the correct answer, the teacher generates a detailed chain-of-thought explanation that highlights the reasoning behind the solution. In this setup, the teacher’s output serves as the preferred answer, while the student’s original response is treated as the rejected answer for Direct Policy Optimization (DPO).
The second framework, depicted in 2, employs the teacher model to provide more granular feedback by identifying specific errors in the student model’s answer in its response to the given multiple choice question. Through this constructive criticism, the student model is guided toward producing the correct answer independently. In this case, the final answer generated by the student is considered the preferred response, while its initial attempt functions as the rejected one for DPO. Although the second framework is considerably more time-consuming due to the iterative feedback process, it can significantly enhance the student model’s learning because DPO leverages preferred-rejected answer pairs generated solely by the student itself.
Due to hardware limitations and the time efficiency of Framework 1, we trained the baseline model, aya-expanse-8b, using Framework 1 for 95% of the data and Framework 2 for the remaining 5%. This approach allowed us to efficiently train the model using Direct Preference Optimization (DPO). In total, the training involved 11,000 preferred–rejected answer pairs, containing approximately 2 million tokens in the preferred answers and 2.5 million tokens in the rejected ones.
Figure 1: Framework 1 Block Diagram
Figure 2: Framework 2 Block Diagram
V Results
In this section, we compare the newly developed gaokerena-R model with its predecessor, gaokerena-V. While gaokerena-V was trained on a large medical corpus and demonstrates strong factual knowledge and retrieval capabilities, gaokerena-R was specifically designed to enhance medical reasoning. Owing to its reasoning-focused training pipeline, gaokerena-R was trained on a substantially smaller dataset, resulting in slightly reduced coverage of general medical knowledge. However, this trade-off enabled it to develop deeper reasoning competence, allowing it to perform better on tasks requiring multi-step inference and logical consistency.
In the final evaluation, we compared the performance of gaokerena-V under direct (straight) prompting with that of gaokerena-R when provided with Chain-of-Thought (CoT) prompts. The results highlight that gaokerena-R, despite its smaller scale and limited training data, achieves superior reasoning performance through structured reasoning guidance, demonstrating the effectiveness of reasoning-centered optimization over pure data scaling.
V-A Medical Reasoning Capabillities
To assess the medical reasoning capabilities of the models, we performed evaluations using the FA_MED_MMLU dataset  [^2] Available at huggingface.co/datasets/gaokerena/FA_MED_MMLU. Both gaokerena-V and gaokerena-R were prompted  [^3] Evaluation prompts are available at github.com/Mehrdadghassabi/Gaokerena-R/blob/main/evaluations/zeroshot-COT/kopp/gaokerena-r1.0/Untitled2.ipynb to generate Chain-of-Thought (CoT) reasoning traces with a temperature setting of 1.0 For each question, five independent samples were generated, and a majority-voting mechanism was applied to determine the final answer: if three or more of the five generations selected the same option, that option was chosen as the final prediction; otherwise, the question was left unanswered to reflect model uncertainty.
This evaluation framework provides a more robust estimation of reasoning consistency and agreement across multiple reasoning trajectories. The results are presented in Table I for the without negative marking setting, and in Table II for the with negative marking setting. These two scoring schemes allow for a fair comparison of the models’ reasoning accuracy under different evaluation criteria.
TABLE I: Chain-of-Thought Prompted Performance Without Negative Marking
gao gao aya-
kerena-R kerena-V expanse-8b
MMLU-
anatomy(fa) 42.22 39.25 40.74
MMLU-
medical-genetics(fa) 50.0 41.0 45.0
MMLU-
college-medicine(fa) 47.97 37.57 48.55
MMLU-
clinical-knowledge(fa) 55.84 46.79 54.71
MMLU-
professional- 44.85 37.13 43.75
medicine(fa)
MMLU-
college-biology(fa) 48.61 36.80 43.75
MMLU(avg) 48.76 40.40 47.10
IBMSEE Sept2023 38.69 29.76 35.71
Number of
parameters 8b 8b 8b
inference time $\approx 5\times 35s$ $\approx 5\times 35s$ $\approx 5\times 35s$
TABLE II: Chain-of-Thought Prompted Performance With Negative Marking
gao gao aya-
kerena-R kerena-V expanse-8b
MMLU-
anatomy(fa) 29.13 27.65 24.93
MMLU-
medical-genetics(fa) 40.0 32.0 33.0
MMLU-
college-medicine(fa) 34.68 25.24 34.48
MMLU-
clinical-knowledge(fa) 44.65 35.59 42.51
MMLU-
professional- 30.39 25.0 30.39
medicine(fa)
MMLU-
college-biology(fa) 36.80 25.0 30.09
MMLU(avg) 36.14 28.57 33.55
IBMSEE Sept2023 24.60 15.87 19.84
Number of
parameters 8b 8b 8b
inference time $\approx 5\times 35s$ $\approx 5\times 35s$ $\approx 5\times 35s$
V-A1 Pass@K
The pass@k metric, originally introduced by B. Brown et al. [25], provides a robust measure for evaluating language models that generate varying outputs for the same input across multiple samples. It quantifies the probability that a model produces at least one correct answer within $k$ independent attempts, thereby offering a more comprehensive assessment of reasoning reliability and sample diversity. The formal definition of the metric is as follows, where $N$ denotes the total number of generated samples and $C_{i}$ represents the number of correct samples for problem $i$:
$\text{pass@k}=\frac{1}{\text{\# of problems}}\sum_{i=1}^{\text{\# of problems}}\left(1-\frac{\binom{N-C_{i}}{k}}{\binom{N}{k}}\right)$ (1)
Following the same experimental setup described earlier, we computed pass@k scores for $k=1,2,3$ using both the FA_MED_MMLU dataset and the IBMSEE (September 2023) dataset. The evaluation included the gaokerena-R, gaokerena-V, and aya-expanse-8b models. The results for the FA_MED_MMLU dataset are illustrated in Figure 3, while the IBMSEE (September 2023) results are shown in Figure 4.
As illustrated in Figures 3 and 4, the gaokerena-R model consistently demonstrated superior performance across nearly all $k$ values and evaluation categories. This improvement indicates that gaokerena-R possesses more stable and coherent reasoning capabilities, producing correct answers more reliably even with a limited number of samples. In contrast, gaokerena-V exhibited relatively weak performance for smaller $k$ values, while its results improved as $k$ increased. This pattern suggests that gaokerena-V is considerably more uncertain when prompted to generate Chain-of-Thought reasoning trajectories, often producing diverse or inconsistent answers across different samples. The results therefore highlight the enhanced reasoning consistency and reliability achieved through gaokerena-R’s targeted reasoning-oriented training approach.
Figure 3: pass@k results on the FA_MED_MMLU dataset using Chain-of-Thought prompting
Figure 4: pass@k results on the IBMSEE (September 2023) dataset using Chain-of-Thought prompting
V-B Medical Knowledge
As mentioned earlier, gaokerena-V possesses a broader base of medical knowledge compared to the newer gaokerena-R model, primarily because it was trained on a significantly larger and more diverse corpus of medical data. This difference becomes evident when both models are evaluated using direct question–answer prompting, rather than being asked to generate explicit Chain-of-Thought reasoning trajectories. Under this direct prompting setting, gaokerena-V demonstrates superior performance, reflecting its stronger memorization and factual recall abilities derived from large-scale medical pretraining. The comparative results for both models on the FA_MED_MMLU and IBMSEE (September 2023) datasets are presented in Table III, which clearly highlights gaokerena-V’s advantage in knowledge-intensive evaluations.
TABLE III: Straight Prompted Performance
gao gao aya-
kerena-R kerena-V expanse-8b
MMLU-
anatomy(fa) 41.48 48.14 40.74
MMLU-
medical-genetics(fa) 49.0 53.0 49.0
MMLU-
college-medicine(fa) 46.24 43.93 44.51
MMLU-
clinical-knowledge(fa) 52.45 55.47 52.07
MMLU-
professional- 41.91 47.05 45.58
medicine(fa)
MMLU-
college-biology(fa) 44.44 47.22 45.14
MMLU(avg) 46.28 49.31 46.64
IBMSEE Sept2023 35.11 38.69 34.52
Number of
parameters 8b 8b 8b
inference time $\approx 10s$ $\approx 10s$ $\approx 10s$
V-C Final Evaluation
As previously discussed, gaokerena-V performs better when prompted directly, whereas gaokerena-R excels when prompted with a Chain-of-Thought (CoT) format that encourages step-by-step reasoning before producing an answer. One key advantage of CoT prompting—particularly when combined with multiple sampling and majority voting—is that it provides a natural measure of model certainty. If the generated responses converge on the same option across samples, the model can be considered confident; conversely, a wide dispersion among responses indicates uncertainty. This approach offers a practical alternative to directly querying the model’s self-assessed confidence [26] , a capability that smaller models generally lack due to their limited self-awareness and introspective reasoning abilities.
In cases where gaokerena-R exhibits uncertainty (i.e., when multiple distinct answers are produced), we employ the baseline model, aya-expanse-8b, as an auxiliary verifier. The outputs from gaokerena-R are presented to aya-expanse-8b, which is tasked with selecting the option containing the least incorrect or inconsistent information. This hybrid evaluation framework enables us to combine the reasoning strengths of gaokerena-R with the broader knowledge coverage of aya-expanse-8b.
Accordingly, we compare two configurations, as summarized in Table IV. In the first configuration , gaokerena-V is evaluated using direct prompting without any reasoning guidance. In the second configuration , gaokerena-R is prompted with a Chain-of-Thought (CoT) format, and in cases of uncertainty, its generated answers are verified by aya-expanse-8b to identify the most reliable response.
TABLE IV: Evaluation of two different configurations
gaokerena-R gaokerena-V
aya-expanse-8b
MMLU-
anatomy(fa) 47.40 48.14
MMLU-
medical-genetics(fa) 56.0 53.0
MMLU-
college-medicine(fa) 50.28 43.93
MMLU-
clinical-knowledge(fa) 58.86 55.47
MMLU-
professional- 48.89 47.05
medicine(fa)
MMLU-
college-biology(fa) 54.86 47.22
MMLU(avg) 52.98 49.31
IBMSEE Sept2023 46.42 38.69
prompt COT for the main model Straight
Straight for the verifier
inference time $\approx 5\times 35+8s$ $\approx 10s$

### v2

@@
 I Introduction
-In 2019, at the NeurIPS conference, Yoshua Bengio highlighted a critical limitation of current deep learning systems: their inability to perform tasks requiring robust reasoning skills [1].
-While these systems excel at intuitive, perception-driven tasks, they struggle with reasoning-based challenges that demand deeper cognitive processing.
-This observation echoes Daniel Kahneman’s distinction between “fast” and “slow” thinking in cognitive science [2].
-Fast, intuitive thinking aligns with the strengths of current AI systems, whereas slow, deliberate reasoning remains a significant weakness.
-One proposed solution to this deficiency is the revival of symbolic AI methods through neurosymbolic approaches.
-However, Bengio cautioned against such methods due to scalability issues.
-Instead, he urged the development of new architectures capable of out-of-distribution generalization—a core requirement for genuine reasoning.
-The Transformer architecture [3], which underpins most modern language models, suffers from the same deficiency.
-Transformers often make “stupid mistakes” when faced with reasoning-intensive problems, prompting researchers to enlarge models and datasets in an attempt to achieve better results across tasks.
-This approach arguably simulates System 2 reasoning within a fundamentally System 1 framework.
-This limitation becomes even more pronounced in smaller language models, where restricted capacity and limited data exacerbate reasoning challenges.
-In domains such as medicine—where reasoning, interpretation, and decision-making are essential—these weaknesses become especially consequential.
-Recent efforts (including ours) to enhance reasoning have largely focused on prompting or fine-tuning methods that enable models to imitate reasoning behavior rather than truly engage in it—a phenomenon we refer to as the “illusion of thinking” [4].
-This means that such approaches provide only an approximation of genuine cognitive reasoning.
-Nevertheless, in the absence of architectures with intrinsic reasoning capabilities, these efforts represent necessary steps toward developing AI systems with more robust reasoning abilities.
-The present work aims to address this challenge by enhancing the reasoning abilities of a small Persian language model, aya-expanse-8b [5].
-We first prompt the model to generate Chain-of-Thought (CoT) trajectories for multiple-choice medical questions.
-These responses are then reviewed and corrected using DeepSeek-R, a model adept at reasoning, to rectify errors.
-This process constitutes Reinforcement Learning with AI Feedback (RLAIF) [6].
-By pairing incorrect and corrected CoT trajectories, we subsequently apply Direct Preference Optimization (DPO) [7] to enhance the reasoning capabilities of the baseline model.
-We call this improved model gaokerena-R.
-While its predecessor, gaokerena-V [8], demonstrated superior medical knowledge in straightforward prompting tasks, gaokerena-R outperformed it when evaluated using Chain-of-Thought (CoT) prompting [9].
-Notably, gaokerena-R was trained on a much smaller dataset, focusing on reasoning enhancement rather than data scale, whereas gaokerena-V relied on extensive training on a large medical corpus and the MF3QA dataset [8].
-Our findings highlight an important insight: in low-resource domains such as Persian medical AI, strengthening reasoning capabilities can be more effective than training on vast datasets.
-This underscores the importance of developing reasoning-oriented approaches, particularly in languages and fields with limited data availability.
-By improving reasoning abilities in smaller models, we can make meaningful progress toward practical, resource-efficient AI systems.
+Transformer-based language models [1] excel at fast, intuitive tasks—such as pattern matching, retrieval, and surface-level text generation—mirroring Kahneman’s concept of “fast thinking” [2].
+However, they struggle with deliberate, multi-step reasoning tasks that require “slow thinking,” particularly in specialized domains such as medicine, where diagnostic accuracy depends on logical inference, evidence evaluation, and error correction.
+This limitation is even more severe in low-resource languages such as Persian, where both high-quality data and compute are scarce.
+In 2019, Yoshua Bengio warned that deep learning systems lack true reasoning capacity and called for architectures that support out-of-distribution generalization [3].
+While the transformer architecture was revolutionary, its success has largely come from scaling: larger models and larger datasets yield better performance.
+Yet, despite current large language models having trillions of parameters and being trained on tens of trillions of tokens, they still make surprisingly simple reasoning errors.
+They may even produce inconsistent answers when asked the same question directly or via chain-of-thought (CoT) prompting [4].
+This deficiency becomes even more pronounced in small medical Persian language models, which have far fewer parameters and much less training data.
+Recent advances have attempted to improve the reasoning ability of language models—enhancing their performance on “slow thinking” tasks—but these methods often rely on large, well-curated datasets available only for high-resource languages such as English or Chinese.
+In contrast, Persian lacks sufficient high-quality medical datasets.
+To address this gap, we propose a new framework to improve the reasoning capabilities of Persian medical language models under limited data availability.
+In our proposed method, we machine-translated an English multiple-choice medical question answering dataset into Persian and applied reinforcement learning from AI feedback (RLAIF) [5] and direct preference optimization (DPO) [6] to enhance the reasoning ability of a baseline Persian medical model.
+The resulting model is named gaokerena-R.
+[^1] All of our work is open-source and available at github.com/Mehrdadghassabi/Gaokerena-R
+Our prior work, gaokerena-V [7], fine-tuned a Persian medical language model using approximately 57 million tokens (including a portion of a medical corpus and a dataset) via supervised fine-tuning (SFT).
+Although gaokerena-V demonstrated strong medical knowledge, gaokerena-R outperforms it when given chain-of-thought prompts—despite being trained on less medical data.
+Since both models share the same baseline model, aya-expanse-8b [8], we hypothesize that enhancing reasoning skills is more beneficial than scaling data for small, low-resource medical language models.
+In summary, our contributions are as follows:
+  1. 1.
+    An efficient RLAIF+DPO framework that generates high-signal CoT preference pairs using a teacher–student loop.
+  2. 2.
+    gaokerena-R  [^2] Available at huggingface.co/gaokerena/gaokerena-r1.0 , an 8b-parameter medical model demonstrating that reasoning-focused training outperforms data scaling in low-resource medical NLP.
+  3. 3.
+    A machine-translated Persian medical multiple-choice question answering dataset designed for reasoning-focused model training.
 II Related Work
 To the best of our knowledge, no prior work has focused on developing Persian medical reasoning language models.
-Existing Persian medical models, including our previous gaokerena-V, mainly address knowledge representation and general language understanding with limited attention to reasoning.
+Existing Persian medical language models, including gaokerena-V, mainly address knowledge representation and general language understanding with limited attention to reasoning.
 L.
 Pan et al.
-reviewed several valuable studies on enhancing reasoning capabilities in language models.
-[10] Accordingly, we focus here on English-language works that offer methodological insights into generating and integrating reasoning data for improving medical language models.
+reviewed several valuable studies on enhancing reasoning capabilities in language models [9].
+Accordingly, we focus here on English-language works that offer methodological insights into generating and integrating reasoning data for improving medical language models.
 II-A Related Work In Medical Domain
@@
 This work demonstrates that reasoning ability can be efficiently transferred across models of different scales through targeted fine-tuning on reasoning-oriented synthetic data.
-III Data
-We required a Persian medical multiple-choice question answering dataset to evaluate our proposed methods.
-In the absence of such a dataset, we used DeepSeek-V3 [20] , a cost-effective large language model, to translate a small portion of the MedMCQA dataset from English to Persian.
-We randomly selected these questions from MedMCQA [21] to maintain topic diversity in the resulting dataset.
-To verify the quality of the translations, we prompted  [^1] Prompts are available at github.com/Mehrdadghassabi/Gaokerena-R/blob/main/dataset/judgement.ipynb two referees (grok-3-mini [22] and gpt-4.1-mini [23] ) to evaluate each translation.
-A translation was considered verified only if the both referees assigned it a score of 5 out of 5.
-This process resulted in approximately 11,000 Persian medical multiple-choice questions.
-III-A PersianMedQA
-Shortly after completing our translation process, a new Persian medical multiple-choice question answering dataset called PersianMedQA was introduced [24] .
-We randomly selected a small portion of its training data and appended it to our dataset to include additional data from a different source.
-IV Proposed Methods
-We proposed two methods.
-The first framework, depicted in 1, utilizes a teacher model ( we used DeepSeek-R), which is itself a reasoning language model, to correct the student model’s medical reasoning mistakes in its response to the given multiple choice question.
+III Proposed Methods
+Two methods has been porposed.
+the first one, depicted in Figure 1, utilizes a teacher model (we used DeepSeek-R), which is itself a reasoning language model, to correct the student model’s medical reasoning mistakes in its response to the given multiple choice question.
 (if the student model chose the correct option we move to the next question in our dataset) When provided with the correct answer, the teacher generates a detailed chain-of-thought explanation that highlights the reasoning behind the solution.
 In this setup, the teacher’s output serves as the preferred answer, while the student’s original response is treated as the rejected answer for Direct Policy Optimization (DPO).
-The second framework, depicted in 2, employs the teacher model to provide more granular feedback by identifying specific errors in the student model’s answer in its response to the given multiple choice question.
-Through this constructive criticism, the student model is guided toward producing the correct answer independently.
-In this case, the final answer generated by the student is considered the preferred response, while its initial attempt functions as the rejected one for DPO.
-Although the second framework is considerably more time-consuming due to the iterative feedback process, it can significantly enhance the student model’s learning because DPO leverages preferred-rejected answer pairs generated solely by the student itself.
+The second method, illustrated in Figure 2, employs the teacher model to provide more granular feedback by identifying specific errors in the student model’s answer to a given multiple-choice question.
+Upon receiving this feedback, the student model is prompted to correct its response precisely from the point where the error occurred.
+Through this iterative process, the student gradually refines its reasoning trajectory and converges on the correct answer.
+This constructive feedback allows the student model to develop reasoning independence, as it learns to correct its own mistakes rather than simply mimicking the teacher’s response.In this setup, the final answer generated by the student is considered the preferred response, while its initial attempt serves as the rejected one for DPO.
+Although this second framework is more time-consuming due to the iterative feedback process, it simplifies DPO training, since both the rejected and preferred answers are generated entirely by the student model itself as sequences of its own tokens.
 Due to hardware limitations and the time efficiency of Framework 1, we trained the baseline model, aya-expanse-8b, using Framework 1 for 95% of the data and Framework 2 for the remaining 5%.
@@
 In total, the training involved 11,000 preferred–rejected answer pairs, containing approximately 2 million tokens in the preferred answers and 2.5 million tokens in the rejected ones.
-Figure 1: Framework 1 Block Diagram
-Figure 2: Framework 2 Block Diagram
-V Results
+Figure 1: Method 1 Block Diagram
+Figure 2: Method 2 Block Diagram
+IV Data
+We required a Persian medical multiple-choice question answering dataset to implement our proposed method.
+In the absence of such a dataset at the time, we used DeepSeek-V3 [19], a cost-effective large language model, to translate a subset of the MedMCQA dataset from English to Persian.
+To maintain topic diversity, the questions were randomly selected from MedMCQA [20].
+To verify the quality of the translations, we prompted  [^3] Prompts are available at github.com/Mehrdadghassabi/Gaokerena-R/blob/main/dataset/judgement.ipynb two referees—grok-3-mini [21] and gpt-4.1-mini [22]—to evaluate each translation.
+A translation was considered verified only if both referees assigned it a score of 5 out of 5.
+This process resulted in approximately 18,000 verified Persian medical multiple-choice questions.
+IV-A PersianMedQA
+Shortly after completing our translation process, a new Persian medical multiple-choice question answering dataset, called PersianMedQA, was introduced [23].
+We randomly selected 1,000 questions from its training set and appended them to our dataset to incorporate additional data from a different source.
+The resulting dataset contained 19,000 Persian medical multiple-choice questions in total.
+V Carbon Footprint
+The carbon footprint of our DPO fine-tuning process was estimated based on the hardware configuration and total runtime.
+The procedure ran for a combined total of 1 hour on an NVIDIA H100 PCIe 80 GB GPU, with approximately 43 GB of VRAM utilized during training.
+The training loss curve is shown in Figure 3.
+Assuming an average power consumption of 350 watts per GPU, the total energy usage was approximately 0.35 kWh.
+Using the average carbon intensity of the Canadian electricity grid, where our server was located (0.086 kilograms of CO2 equivalent per kWh [24]), this corresponds to an estimated emission of 0.0301 kilograms of CO2 equivalent generated during the fine-tuning process.
+Compared to our previous model, gaokerena-V, which emitted 2.66 kilograms of CO2, this represents a substantial reduction in environmental impact.[25]
+Figure 3: Training loss curve
+VI Results
 In this section, we compare the newly developed gaokerena-R model with its predecessor, gaokerena-V.
@@
 The results highlight that gaokerena-R, despite its smaller scale and limited training data, achieves superior reasoning performance through structured reasoning guidance, demonstrating the effectiveness of reasoning-centered optimization over pure data scaling.
-V-A Medical Reasoning Capabillities
-To assess the medical reasoning capabilities of the models, we performed evaluations using the FA_MED_MMLU dataset  [^2] Available at huggingface.co/datasets/gaokerena/FA_MED_MMLU.
-Both gaokerena-V and gaokerena-R were prompted  [^3] Evaluation prompts are available at github.com/Mehrdadghassabi/Gaokerena-R/blob/main/evaluations/zeroshot-COT/kopp/gaokerena-r1.0/Untitled2.ipynb to generate Chain-of-Thought (CoT) reasoning traces with a temperature setting of 1.0 For each question, five independent samples were generated, and a majority-voting mechanism was applied to determine the final answer: if three or more of the five generations selected the same option, that option was chosen as the final prediction; otherwise, the question was left unanswered to reflect model uncertainty.
-This evaluation framework provides a more robust estimation of reasoning consistency and agreement across multiple reasoning trajectories.
-The results are presented in Table I for the without negative marking setting, and in Table II for the with negative marking setting.
+VI-A Medical Reasoning Capabillities
+To assess the medical reasoning capabilities of the models, we prompted them to produce Chain-of-Thought reasoning trajectories with a temperature setting of 1.0.
+Since all models can produce different answers for the same question  [^4] Evaluation prompts are available at github.com/Mehrdadghassabi/Gaokerena-R/blob/main/evaluations/zeroshot-COT/kopp/gaokerena-r1.0/Untitled2.ipynb, we evaluated them using two metrics on the FA_MED_MMLU  [^5] Available at huggingface.co/datasets/gaokerena/FA_MED_MMLU and IBMSEE (September 2023)  [^6] Available at huggingface.co/datasets/gaokerena/KOPP datasets.
+VI-A1 Accuracy
+The first metric is accuracy for each question in the datasets.
+Five independent samples were generated per question, and a majority-voting mechanism was applied: if three or more of the five generations selected the same option, that option was chosen as the final prediction; otherwise, the question was left unanswered to reflect model uncertainty.
+This framework provides a robust estimate of reasoning consistency across multiple trajectories.
+Results are presented in Table I (without negative marking) and Table II (with negative marking), where, in the negative marking setting, each incorrectly answered question is assigned a score of -0.33.
 These two scoring schemes allow for a fair comparison of the models’ reasoning accuracy under different evaluation criteria.
@@
 inference time $\approx 5\times 35s$ $\approx 5\times 35s$ $\approx 5\times 35s$
-V-A1 Pass@K
+VI-A2 Pass@K
 The pass@k metric, originally introduced by B.
@@
 It quantifies the probability that a model produces at least one correct answer within $k$ independent attempts, thereby offering a more comprehensive assessment of reasoning reliability and sample diversity.
-The formal definition of the metric is as follows, where $N$ denotes the total number of generated samples and $C_{i}$ represents the number of correct samples for problem $i$:
+The formal definition of the metric is given in Formula 1 , where $N$ denotes the total number of generated samples and $C_{i}$ represents the number of correct samples for problem $i$.
 $\text{pass@k}=\frac{1}{\text{\# of problems}}\sum_{i=1}^{\text{\# of problems}}\left(1-\frac{\binom{N-C_{i}}{k}}{\binom{N}{k}}\right)$ (1)
@@
 The results for the FA_MED_MMLU dataset are illustrated in Figure 4, while the IBMSEE (September 2023) results are shown in Figure 5.
-As illustrated in Figures 3 and 4, the gaokerena-R model consistently demonstrated superior performance across nearly all $k$ values and evaluation categories.
+As illustrated in Figures 4 and 5, the gaokerena-R model consistently demonstrated superior performance across nearly all $k$ values and evaluation categories.
 This improvement indicates that gaokerena-R possesses more stable and coherent reasoning capabilities, producing correct answers more reliably even with a limited number of samples.
@@
 Figure 5: Pass@k results on the IBMSEE (September 2023) dataset using Chain-of-Thought prompting
-V-B Medical Knowledge
+VI-B Medical Knowledge
 As mentioned earlier, gaokerena-V possesses a broader base of medical knowledge compared to the newer gaokerena-R model, primarily because it was trained on a significantly larger and more diverse corpus of medical data.
 This difference becomes evident when both models are evaluated using direct question–answer prompting, rather than being asked to generate explicit chain-of-thought reasoning trajectories.
-Under this direct prompting setting, gaokerena-V demonstrates superior performance, reflecting its stronger memorization and factual recall abilities derived from large-scale medical pretraining.
-The comparative results for both models on the FA_MED_MMLU and IBMSEE (September 2023) datasets are presented in Table III, which clearly highlights gaokerena-V’s advantage in knowledge-intensive evaluations.
+As shown in Table III, under this direct prompting setting, gaokerena-V demonstrates superior performance, reflecting its stronger memorization and factual recall abilities derived from large-scale medical pretraining.
+Interestingly, the approximately equal performance of gaokerena-R and its baseline aya-expanse-8b in this same direct prompting setup indicates that gaokerena-R’s superior results in the chain-of-thought setting stem primarily from its enhanced reasoning skills rather than from increased medical knowledge.
 TABLE III: Straight Prompted Performance
@@
 inference time $\approx 10s$ $\approx 10s$ $\approx 10s$
-V-C Final Evaluation
+VI-C Final Evaluation
 As previously discussed, gaokerena-V performs better when prompted directly, whereas gaokerena-R excels when prompted with a Chain-of-Thought (CoT) format that encourages step-by-step reasoning before producing an answer.
@@
 inference time $\approx 5\times 35+8s$ $\approx 10s$
+VII Future Research
+The results show that gaokerena-V performs better with direct prompts, while gaokerena-R excels with chain-of-thought (CoT) prompts.
+This indicates that both models are prompt-dependent, as their performance varies with the prompting style.
+Future research should aim to develop prompt-invariant medical language models that integrate strong reasoning skills and medical knowledge, achieving superior performance regardless of the prompt format.
+Achieving prompt invariance would represent an important step toward more reliable and generalizable small Persian medical language models.

### v3

@@
 To accomplish this, the authors applied the Monte Carlo Tree Search (MCTS) [11] algorithm on medical multiple-choice question-answering datasets to synthesize structured reasoning trajectories using a policy model.
-Based on this approach, they constructed three datasets supporting different stages of training: a dataset for Supervised Fine-Tuning (SFT) of the policy model, a rejected–preferred answer dataset for Direct Preference Optimization (DPO) [12] ,and a soft dual-side label dataset for fine-tuning the Process Reward Model (PRM).Finally, the trained Policy Model was used as the primary reasoning policy, while the trained Process Reward Model acted as a decoding guide to refine and evaluate the reasoning process during generation.
+Based on this approach, they constructed three datasets supporting different stages of training: a dataset for Supervised Fine-Tuning (SFT) of the policy model, a rejected–preferred answer dataset for Direct Preference Optimization (DPO) ,and a soft dual-side label dataset for fine-tuning the Process Reward Model (PRM).Finally, the trained Policy Model was used as the primary reasoning policy, while the trained Process Reward Model acted as a decoding guide to refine and evaluate the reasoning process during generation.
 Another significant contribution in this field is MedReason [12].
@@
 Perhaps the most influential recent work in the broader field of AI reasoning is DeepSeek-R [14].
-Building upon the DeepSeek-V3-Base model, the authors introduced a reinforcement learning framework based on Gradient Regularized Policy Optimization (GRPO) [16] to explicitly enhance the model’s reasoning capabilities.
+Building upon the DeepSeek-V3-Base model, the authors introduced a reinforcement learning framework based on Group Relative Policy Optimization (GRPO) [15] to explicitly enhance the model’s reasoning capabilities.
 In this setup, the reinforcement learning process was guided by a reward function specifically designed to evaluate and maximize reasoning performance.
@@
 (if the student model chose the correct option we move to the next question in our dataset) When provided with the correct answer, the teacher generates a detailed chain-of-thought explanation that highlights the reasoning behind the solution.
-In this setup, the teacher’s output serves as the preferred answer, while the student’s original response is treated as the rejected answer for Direct Policy Optimization (DPO).
+In this setup, the teacher’s output serves as the preferred answer, while the student’s original response is treated as the rejected answer for Direct Preference Optimization (DPO).
 The second method, illustrated in Figure 2, employs the teacher model to provide more granular feedback by identifying specific errors in the student model’s answer to a given multiple-choice question.
@@
 The carbon footprint of our DPO fine-tuning process was estimated based on the hardware configuration and total runtime.
-The procedure ran for a combined total of 1 hour on an NVIDIA H100 PCIe 80 GB GPU, with approximately 43 GB of VRAM utilized during training.
+The procedure ran for a combined total of 1 hour  [^4] The reported 1 hour refers only to the training time and does not include the time spent on data generation.
+on an NVIDIA H100 PCIe 80 GB GPU, with approximately 43 GB of VRAM utilized during training.
 The training loss curve is shown in Figure 3.
@@
 Straight for the verifier
-inference time $\approx 5\times 35+8s$ $\approx 10s$
+inference time $\approx 5\times 35+10+8s$ $\approx 10s$
 VII Future Research

### v4

@@
 II-A Related Work In Medical Domain
-One of the notable efforts in developing small-scale medical reasoning language models is MedSSS [10].
-This framework focuses on fine-graining the intermediate reasoning steps to enhance the reasoning ability of medical models.
-To accomplish this, the authors applied the Monte Carlo Tree Search (MCTS) [11] algorithm on medical multiple-choice question-answering datasets to synthesize structured reasoning trajectories using a policy model.
-Based on this approach, they constructed three datasets supporting different stages of training: a dataset for Supervised Fine-Tuning (SFT) of the policy model, a rejected–preferred answer dataset for Direct Preference Optimization (DPO) ,and a soft dual-side label dataset for fine-tuning the Process Reward Model (PRM).Finally, the trained Policy Model was used as the primary reasoning policy, while the trained Process Reward Model acted as a decoding guide to refine and evaluate the reasoning process during generation.
-Another significant contribution in this field is MedReason [12].
-In this work, the authors leveraged a structured medical knowledge graph to transform conventional question–answer (Q&A) pairs into detailed medical reasoning trajectories.
-Each trajectory represented the step-by-step logical pathway connecting a clinical question to its correct answer, grounded in medical knowledge relationships such as symptoms, diagnoses, treatments, and physiological mechanisms.
-By constructing a reasoning-enriched dataset in this manner, the authors were able to fine-tune a baseline language model on these structured reasoning examples.
-This approach led to a marked improvement in the model’s ability to perform reasoning-intensive tasks within the medical domain, demonstrating the effectiveness of incorporating knowledge graph–based reasoning supervision into the training process.
-The findings from MedReason further emphasize the importance of structured reasoning representations as a means of improving the interpretability and analytical depth of medical language models.
-Another important advancement in the development of medical reasoning models is HuatuoGPT-o1 [13].
-In this work, the authors introduced a verification-guided reasoning framework designed to improve the logical consistency and accuracy of generated reasoning trajectories.
-Specifically, they employed a verifier model to assess and guide the policy model during reasoning generation, ensuring that each reasoning path adhered closely to factual correctness and medical plausibility.
-By filtering and refining reasoning trajectories through this verification process, they created a high-quality dataset composed of verified reasoning sequences.
-The authors then leveraged both supervised fine-tuning and reinforcement learning techniques to train their baseline language model using these verified data.
-This dual training approach allowed the model to not only imitate correct reasoning behaviors but also internalize reasoning principles through reward-driven optimization.
-As a result, HuatuoGPT-o1 demonstrated significant improvements in reasoning accuracy and reliability across a variety of medical question-answering and diagnostic tasks, highlighting the potential of verifier-guided learning frameworks in advancing medical reasoning language models.
+The MedSSS framework is a notable, self-evolving system designed to instill robust, long-chain reasoning capabilities into small, deployable medical language models, such as the Llama3.1-8B-Instruct base model [10].
+Its core methodology involves leveraging Monte Carlo Tree Search (MCTS) [11] over diverse medical datasets to construct rule-verifiable reasoning trajectories.
+These generated paths are used for policy refinement via supervised fine-tuning (SFT) and direct preference optimization (DPO), and to train a unique Process Reward Model (PRM).
+This PRM employs a soft dual-sided labeling objective that provides crucial step-level supervision, penalizing reasoning steps that degrade node value to mitigate hallucination and enhance interpretability in complex clinical reasoning.
+The resulting MedSSS system demonstrated state-of-the-art performance across eleven clinical reasoning benchmarks, achieving an average performance gain of +14.12 compared to its base model and significantly surpassing 32B-scale general-purpose reasoning models.
+The MedReason [12] framework is another significant contribution that addresses the scarcity of high-quality medical reasoning data by leveraging a structured medical knowledge graph (KG).
+This approach converts conventional question–answer (Q&A) pairs into a dataset of 32,682 factually grounded reasoning trajectories, serving as structural supervision.
+Fine-tuning models on these paths consistently boosts performance: the framework achieved average accuracy gains of +5.4% for LLaMA 3.1-Instruct-8B and up to 7.7% for DeepSeek-Distill-8B.
+The resulting MedReason-8B model established a new state-of-the-art among 8B-parameter models, notably surpassing the Huatuo-o1-RL-8B by up to 4.2% on the MedBullets clinical benchmark, validating the importance of KG-based reasoning for interpretability and analytical depth.
+The HuatuoGPT-o1 [13] framework is a two-stage, verification-guided system designed to instill complex reasoning in medical LLMs using 40K verifiable medical problems.
+Initially, a verifier guides search strategies to generate verified reasoning trajectories for Supervised Fine-Tuning (SFT), teaching the model to refine its answers.
+Subsequently, Reinforcement Learning (RL) is applied using verifier-based rewards for further enhancement.
+This methodology resulted in an average 8.5-point performance gain for the 8B model on medical benchmarks, with the complex Chain-of-Thought (CoT) strategy alone providing an average 4.3-point boost.
 II-B Related Work In Other Domain
-Perhaps the most influential recent work in the broader field of AI reasoning is DeepSeek-R [14].
-Building upon the DeepSeek-V3-Base model, the authors introduced a reinforcement learning framework based on Group Relative Policy Optimization (GRPO) [15] to explicitly enhance the model’s reasoning capabilities.
-In this setup, the reinforcement learning process was guided by a reward function specifically designed to evaluate and maximize reasoning performance.
-Through this training paradigm, DeepSeek-R demonstrated remarkable improvements in logical reasoning and problem-solving accuracy across various benchmarks.
-However, the application of reinforcement learning also introduced several side effects.
-While reasoning performance improved substantially, the model’s readability and linguistic coherence deteriorated, and instances of language mixing became more frequent.
-To address these issues, the authors incorporated a small amount of cold-start supervised data and adopted a multi-stage training pipeline.
-This additional fine-tuning phase helped restore natural language fluency and readability while retaining the strong reasoning skills acquired through reinforcement learning.
-The resulting model, DeepSeek-R, thus represents a critical step forward in reasoning-oriented AI, demonstrating that reinforcement learning can significantly enhance reasoning ability—provided it is balanced with targeted fine-tuning to maintain linguistic quality.
-Another notable contribution in the area of reasoning enhancement is Thought Preference Optimization (TPO) [16].
-In this work, the authors proposed a preference-based framework for improving reasoning quality in language models.
-Given a question, the model first generates multiple candidate reasoning trajectories.
-These responses are then evaluated by a judge model, which identifies the best and worst samples based on reasoning correctness and coherence.
-The collected best–worst pairs are subsequently used to train the baseline model through Direct Preference Optimization (DPO), encouraging it to prefer higher-quality reasoning paths.
-This approach effectively aligns the model’s reasoning process with human-like evaluative feedback, demonstrating that reasoning quality can be substantially improved through structured preference optimization rather than relying solely on scale or supervised data.
-Another relevant study was conducted by N.
+The DeepSeek-R1 model [14] demonstrated that applying a reinforcement learning framework, specifically the resource-efficient Group Relative Policy Optimization (GRPO) [15] , to the DeepSeek-V3-Base model can effectively enhance complex reasoning capabilities.
+While the pure RL training (DeepSeek-R1-Zero) yielded powerful reasoning behaviors, it simultaneously introduced significant side effects, most notably a deterioration in readability and an increase in language mixing.
+To counteract these issues, the authors implemented a multi-stage curriculum that incorporated a cold-start supervised fine-tuning (SFT) phase.
+This targeted intervention proved essential for restoring the model’s linguistic quality and coherence.
+The resultant model, DeepSeek-R1, leveraged this carefully balanced training pipeline to achieve reasoning performance comparable to state-of-the-art models such as OpenAI-o1-1217, thereby confirming that the successful enhancement of reasoning through RL requires a subsequent, targeted fine-tuning stage to maintain linguistic integrity and coherence.
+A notable contribution in the area of reasoning enhancement is Thought Preference Optimization (TPO), proposed by Wu et al [16].
+as an iterative Reinforcement Learning from AI Feedback (RLAIF) approach.
+This preference-based framework compels the model to generate multiple candidate thought and response pairs, subsequently utilizing a judge model to evaluate only the resulting response quality.
+The collected best–worst pairs are then used to train the baseline model via Direct Preference Optimization (DPO), teaching it to generate high-quality internal thoughts without direct thought supervision.
+Utilizing an 8B parameter model as the base, this method demonstrated substantial performance gains in general instruction following, achieving an impressive 52.5% win rate on AlpacaEval (LC) and 37.3% on Arena-Hard.
+These results represent gains of over 4% compared to the direct response baseline, confirming that structured preference optimization effectively enhances reasoning capabilities, even across non-traditional domains such as marketing and health.
+The study by N.
 Ho et al.
-[17], who explored the transfer of reasoning capabilities from large language models to smaller ones.
-In their approach, a smaller model was fine-tuned using data generated by a larger model that exhibited stronger reasoning performance.
-The larger model produced reasoning trajectories and question–answer pairs that served as high-quality supervision signals for the smaller model.
-Through this distillation process, the smaller model effectively learned reasoning strategies and problem-solving patterns from its larger counterpart, achieving competitive reasoning performance with significantly reduced computational cost.
-This work demonstrates that reasoning ability can be efficiently transferred across models of different scales through targeted fine-tuning on reasoning-oriented synthetic data.
+[17] introduced Fine-tune-CoT, an effective knowledge distillation method that leverages prohibitively large language models (LLMs) to generate Chain-of-Thought (CoT) rationales for the fine-tuning of significantly smaller student models.
+This systematic transfer process enabled student models, which were approximately 25–100x smaller in parameter count, to acquire substantial reasoning capabilities.
+For instance, the 6.7B student model achieved an accuracy of 53.33% on MultiArith using diverse reasoning, demonstrating that complex, high-performance reasoning abilities can be efficiently transferred across a significant reduction in model scale.
 III Proposed Methods
-Two methods has been porposed.
-the first one, depicted in Figure 1, utilizes a teacher model (we used DeepSeek-R), which is itself a reasoning language model, to correct the student model’s medical reasoning mistakes in its response to the given multiple choice question.
-(if the student model chose the correct option we move to the next question in our dataset) When provided with the correct answer, the teacher generates a detailed chain-of-thought explanation that highlights the reasoning behind the solution.
-In this setup, the teacher’s output serves as the preferred answer, while the student’s original response is treated as the rejected answer for Direct Preference Optimization (DPO).
-The second method, illustrated in Figure 2, employs the teacher model to provide more granular feedback by identifying specific errors in the student model’s answer to a given multiple-choice question.
-Upon receiving this feedback, the student model is prompted to correct its response precisely from the point where the error occurred.
-Through this iterative process, the student gradually refines its reasoning trajectory and converges on the correct answer.
-This constructive feedback allows the student model to develop reasoning independence, as it learns to correct its own mistakes rather than simply mimicking the teacher’s response.In this setup, the final answer generated by the student is considered the preferred response, while its initial attempt serves as the rejected one for DPO.
-Although this second framework is more time-consuming due to the iterative feedback process, it simplifies DPO training, since both the rejected and preferred answers are generated entirely by the student model itself as sequences of its own tokens.
-Due to hardware limitations and the time efficiency of Framework 1, we trained the baseline model, aya-expanse-8b, using Framework 1 for 95% of the data and Framework 2 for the remaining 5%.
-This approach allowed us to efficiently train the model using Direct Preference Optimization (DPO).
-In total, the training involved 11,000 preferred–rejected answer pairs, containing approximately 2 million tokens in the preferred answers and 2.5 million tokens in the rejected ones.
+We propose a two-stage framework to enhance the reasoning capabilities of small Persian medical language models.
+First, we employ Reinforcement Learning with AI Feedback (RLAIF) to construct a preference dataset $\mathcal{D}=\{(x,y_{w},y_{l})\}$, where $x$ is a medical question, $y_{w}$ is a preferred reasoning trajectory, and $y_{l}$ is a rejected one.
+Second, we fine-tune the student model $\pi_{\theta}$ using Direct Preference Optimization (DPO).
+To generate high-quality preference pairs, we utilize a teacher-student architecture.
+The student ($\pi_{S}$) is the baseline Persian model, and the teacher ($\pi_{T}$) is a reasoning-capable model (DeepSeek-R).
+We employ two strategies to populate $\mathcal{D}$ based on the student’s performance:
+  1. 1.
+    Teacher Correction for Reasoning Alignment: This strategy accounts for 95% of our training data.
+The student model $\pi_{S}$ generates an initial response $y_{init}$ for a given question $x$.
+If the response is incorrect, $y_{init}$ is designated as the rejected response ($y_{l}$).
+The teacher model $\pi_{T}$ is then prompted with the ground truth answer and instructed to generate a correct, detailed Chain-of-Thought (CoT) explanation.
+This teacher-generated trajectory serves as the preferred response ($y_{w}$).
+The resulting preference pair $(x,y_{w},y_{l})$ explicitly contrasts flawed student reasoning with expert teacher reasoning.
+This approach is depicted in Figure 1.
+  2. 2.
+    Teacher-Guided Self-Correction: For the remaining 5% of the data, we employ an iterative feedback loop to encourage intrinsic error correction.
+This limited scale was primarily due to hardware constraints.
+If the student’s initial response $y_{init}$ is incorrect, the teacher generates a textual critique $c$ that identifies the error without revealing the correct option.
+The student is then prompted with $c$ to self-correct and generate a new response, $y_{retry}$.
+If $y_{retry}$ is correct, it is designated as the preferred response ($y_{w}$), and the original incorrect response $y_{init}$ remains the rejected response ($y_{l}$).
+This method ensures that the final $y_{w}$ is generated by the student model’s own distribution.
+The full process is detailed in Figure 2.
+We optimize the student policy $\pi_{\theta}$ to satisfy the generated preferences using Direct Preference Optimization (DPO).
+DPO is an alignment method that avoids the need for a separate, expensive reward model.
+Instead, it optimizes the policy directly by minimizing the negative log-likelihood of the preferred response relative to the reference model $\pi_{ref}$ (the frozen aya-expanse-8b).
+The objective function is defined as:
+$\begin{split}\mathcal{L}_{DPO}(\pi_{\theta};\pi_{ref})=&-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\\ &\hskip-20.00003pt\Biggl[\log\sigma\Biggl(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)}\Biggr)\Biggr]\end{split}$
+Where $\sigma$ is the sigmoid function and $\beta$ is a hyperparameter that controls the extent of the policy’s deviation from the reference model.
+This single objective implicitly maximizes the probability of the valid reasoning trajectory $y_{w}$ while suppressing the probability of the flawed reasoning path $y_{l}$.
+The final dataset comprised approximately 11,000 pairs, yielding about 2 million preferred tokens and 2.5 million rejected tokens.
 Figure 1: Method 1 Block Diagram
@@
 This process resulted in approximately 18,000 verified Persian medical multiple-choice questions.
-IV-A PersianMedQA
-Shortly after completing our translation process, a new Persian medical multiple-choice question answering dataset, called PersianMedQA, was introduced [22].
-We randomly selected 1,000 questions from its training set and appended them to our dataset to incorporate additional data from a different source.
-The resulting dataset contained 19,000 Persian medical multiple-choice questions in total.
 V Carbon Footprint
@@
 The results highlight that gaokerena-R, despite its smaller scale and limited training data, achieves superior reasoning performance through structured reasoning guidance, demonstrating the effectiveness of reasoning-centered optimization over pure data scaling.
+[^5] You can see some examples that show the enhancements in the Gaokerena-R model at github.com/Mehrdadghassabi/Gaokerena-R/tree/main/assets/examples
 VI-A Medical Reasoning Capabillities
@@
 As previously discussed, gaokerena-V performs better when prompted directly, whereas gaokerena-R excels when prompted with a Chain-of-Thought (CoT) format that encourages step-by-step reasoning before producing an answer.
+This observation indicates a high degree of prompt dependence, a key characteristic of both models in their current state.
 One key advantage of CoT prompting—particularly when combined with multiple sampling and majority voting—is that it provides a natural measure of model certainty.
@@
 The results show that gaokerena-V performs better with direct prompts, while gaokerena-R excels with chain-of-thought (CoT) prompts.
-This indicates that both models are prompt-dependent, as their performance varies with the prompting style.
-Future research should aim to develop prompt-invariant medical language models that integrate strong reasoning skills and medical knowledge, achieving superior performance regardless of the prompt format.
+This indicates that both models are prompt-dependent, and future work must address this critical limitation.
+Specifically, future research should aim to develop prompt-invariant medical language models that integrate strong reasoning skills and medical knowledge, achieving superior performance regardless of the prompt format.
+This involves resolving the inherent trade-off between the superior factual recall of a data-scaled model (like gaokerena-V) and the enhanced logical consistency of a reasoning-focused model (like gaokerena-R).
 Achieving prompt invariance would represent an important step toward more reliable and generalizable small Persian medical language models.